# FORECASTOR – II. Simulating Galaxy Surveys with the Cosmological Advanced Survey Telescope for Optical and UV Research

### v1

@@ 5.1 Source Extraction
For each of the extracted sources, we find the corresponding input source by matching the input and extracted catalogue source positions. To reduce computational expense, we only include input galaxies with magnitude mag in our search list, as from our point source analysis we find that no galaxies will be detected beyond this threshold (see below). This matching allows us to determine which of the input galaxies are detected, and thus analyse the completeness as functions of various input parameters. This also allows us to compare the input and measured properties to determine the accuracy of our CASTOR measurements.
@@ 5.2 Magnitude Completeness
Wide | 24.0 | 22.5 | 24.6 | 23.6 | 24.9 |
Deep | 27.3 | 26.3 | 27.3 | 26.9 | 27.1 |
Ultra-Deep | 28.9 | 28.0 | 28.6 | 28.3 | 28.2 |
@@
Wide | 28.2 | 27.3 | 28.0 | 27.6 | 27.9 |
Deep | 30.0 | 29.2 | 29.7 | 29.4 | 29.5 |
Ultra-Deep | 31.0 | 30.4 | 30.9 | 30.6 | 30.7 |
@@ 5.2.2 Galaxies
To investigate the success of CASTOR in detecting extended galaxies, we consider mock images of six CASTOR pointings, taken from independent, non-overlapping regions of the input 2 deg catalogue. This allows us to account for some cosmic variance. We extract and then match all detected galaxies within each of these six images, for each filter and survey exposure time.
@@
The Ultra-Deep survey with the largest exposure times clearly detects the most galaxies, with the largest detection fractions out to the faintest magnitudes. As shown in Table 2, the 75% completeness limits for this Ultra-Deep survey are 28.0–28.9 mag, with galaxies as faint as 31 mag detected. The filters are well matched in depth within this Ultra-Deep survey. The Deep survey will reach around 1–1.7 magnitudes brighter than the Ultra-Deep survey, with a factor of 10 less exposure time.
The significantly shallower Wide survey will be 75% complete between 22.5–24.9 mag, approximately 5 mag fainter than the Ultra-Deep survey, and will reach depths of up to 28 mag. In this survey the decreased sensitivity of the UV-split and u-split filters is clear, resulting in a decreased completeness limit of up to 1.5 mag compared to their un-split counterparts. These narrower filters will play a key role in understanding the properties of the sources detected in CASTOR, providing further data points for SED fitting, as well as being combined with the un-split filter images to create even deeper detection images, however in general the UV, u and g filters will be the optimal filters for source detection in the Wide survey.
To further compare the performance of each survey, we list the number of galaxies detected in each CASTOR FOV in Table 3, as well as an estimation of the total number of galaxies in the full surveys by extrapolating this value across the survey areas. In one pointing, the Ultra-Deep survey is expected to detect an order of magnitude more galaxies than the Wide survey, of order 100,000 galaxies compared to . However, with its large survey area of 2227 deg, this will correspond to detected galaxies across the whole Wide survey. The much deeper Ultra-Deep survey’s small volume of only 4 CASTOR pointings (1 deg), will detect galaxies, albeit down to much fainter magnitudes. The intermediate Deep survey, covering 83 deg, will detect galaxies.
@@ 5.3 Completeness relative to galaxy properties
We now investigate the detectability of galaxies in each CASTOR survey relative to the intrinsic galaxy properties from the input catalogue, for galaxies brighter than mag. In Figure 6 we show the fraction of input galaxies detected as a function of stellar mass for each survey. This is shown for each of the filters, and split into redshift bins. The detection fraction generally increases with galaxy stellar mass, with a plateau at large mass depending on the filter and exposure time. Generally, we expect the Wide survey to detect very few galaxies with . The Deep survey may detect galaxies down to , while the Ultra-Deep survey may detect of galaxies down to .
In Figure 7 we show the equivalent figure of the fraction of galaxies detected as a function of galaxy disk radius. Galaxies with very small disk sizes of kpc have much lower rates of successful detection, although there is otherwise minor variation of detection fraction with disk size. The average detection fraction of disks with kpc depends strongly on exposure time and filter, with the Wide survey having detection fractions of , up to with the Ultra-Deep survey.
@@ 5.4 Magnitude recovery
Finally we consider how the input and recovered magnitudes compare.
In Figure 9 we show the extracted ISO magnitudes from Source Extractor, compared with the input magnitude. We note that throughout this paper we have otherwise considered only the input theoretical magnitude. We find that the input and extracted magnitudes are in good agreement, for all filters and each of the three surveys.
We find that in the Wide survey, 95% of all galaxies have an extracted magnitude within 9.2%, 10.6%, 7.9%, 8.9% and 7.5% of the input magnitude, for the UV, UV-split, u, u-split, and g filters respectively. These are slightly improved in the Deep and Ultra-Deep surveys, to 6.8%, 7.1%, 5.8%, 6.1% and 5.5% for Deep, and 6.0%, 6.2%, 6.1%, 5.7% and 5.9% for Ultra-Deep. The low number of galaxies with larger discrepancies between the input and output magnitudes are very faint.
Overall, the CASTOR surveys can accurately measure galaxy magnitudes.
@@ 6.1 Synergy with Euclid and Roman
CASTOR’s wide-field survey capabilities and UV coverage will play an indispensable role in the new era of wide-field near-infrared extragalactic surveys, alongside Roman (Spergel et al., 2013, 2015) and Euclid (Laureijs et al., 2011; Racca et al., 2016). These space-based observatories have fields of view of the order of hundred times bigger than Hubble’s (0.54 deg for Euclid and 0.28 deg for Roman) and possess angular resolutions comparable to that of Hubble (with a mirror diameter of 1.2 m for Euclid and 2.4 m for Roman), similar to those of CASTOR (0.25 deg and m). Roman and Euclid focus on the near-infrared and optical, covering wavelengths up to m. However, as these will provide no wavelength coverage below Å, CASTOR will play a key role with its wavelength coverage that extends beyond the optical to the near UV. Together, these three observatories are set to transform galaxy and cosmological surveys. Harnessing the synergy between CASTOR and other observatories will be the key to maximizing the scientific return across all observatories.
@@ 7 Conclusions
The Cosmological Advanced Survey Telescope for Optical and UV Research (CASTOR) is a planned Canadian-led space telescope designed to study the Universe in the UV and optical. A core component of the planning for CASTOR is the preparation of a suite of mission planning and science simulation tools, the Finding Optical Requirements and Exposure times for CASTOR (FORECASTOR) project. In this work we present the CASTOR image simulator, a generator which simulates extragalactic imaging for CASTOR observation planning. This generator is built upon the framework of the GalSim Python package, and creates images containing sources from simulated light-cones from the Santa Cruz Semi-Analytic Model.
@@
Three main extragalactic surveys planned for the mission are the Wide, Deep, and Ultra-Deep Legacy Surveys, spanning 2227, 83 and 1 deg, respectively. These surveys will reach total exposure times of 1000, 18,000 and 180,000s respectively in the UV, UV-split, u and u-split filters, with twice the exposure time in the g filter. In this paper we present mock images of each of these surveys, and investigate the detectability of galaxies at these planned exposure times. These images are publicly available222Mock CASTOR survey images generated in this work are available at https://doi.org/10.11570/24.0003, to enable further investigations of the extragalactic science achievable with CASTOR. The resolution of CASTOR will result in crisp UV images, which will significantly out-perform GALEX with its that resulted in significant galaxy blending. This is also a higher resolution than for Rubin, which is limited to due to atmospheric seeing.
@@
We extract sources from the images and compare those to the input source catalogue, in order to estimate the completeness of each CASTOR survey. For point sources, we predict that the Wide survey will be 75% complete down to 26.2–27.2 mag, with the Deep survey extending to 28.2–28.8 mag, and the Ultra-Deep survey reaching to 29.5–30.2 mag. The five filters, UV, UV-split, u, u-split, and g, are generally well matched in depth for each survey. For galaxies, we expect that the faintest detectable galaxy discovered in the Wide survey will have mag, with Deep reaching to mag and Ultra-Deep discovering some faint sources out to mag. The recovered magnitudes are in good agreement with those from the input catalogue. For the completeness relative to stellar mass, we expect the Wide survey to detect very few galaxies with . The Deep survey may detect galaxies down to , while the Ultra-Deep survey may detect of galaxies down to . Galaxies with disk radii of kpc will be the most difficult to detect.
@@
For a single 0.25 deg pointing, the Wide survey is expected to detect of order galaxies, with galaxies seen in one pointing of the Deep and Ultra-Deep surveys. Extending these counts to the full survey areas, we expect the Wide survey to detect galaxies, with galaxies in the Deep survey and in the Ultra-Deep survey.

### v2

@@ 5.1 Source Extraction
From each of the input source positions, we find the nearest extracted source. If the matched source is within , we consider the input source as having been successfully detected. We do not include sources in our detected or input number counts that have multiple matches within , for simplicity. To reduce computational expense, we only include input galaxies with magnitude mag in our search list, as from our point source analysis we find that no galaxies will be detected beyond this threshold (see below). This matching allows us to determine which of the input galaxies are detected, and thus analyse the completeness as functions of various input parameters. This also allows us to compare the input and measured properties to determine the accuracy of our CASTOR measurements.
@@ 5.2 Magnitude Completeness
Wide | 24.8 | 22.8 | 24.9 | 24.0 | 25.0 |
Deep | 27.6 | 26.7 | 27.6 | 27.2 | 27.3 |
Ultra-Deep | 29.2 | 28.3 | 28.9 | 28.7 | 28.7 |
@@
Wide | 27.5 | 26.6 | 27.4 | 26.9 | 27.2 |
Deep | 29.4 | 28.6 | 29.2 | 28.9 | 29.0 |
Ultra-Deep | 30.7 | 29.9 | 30.5 | 30.2 | 30.3 |
@@ 5.2.2 Galaxies
To investigate the success of CASTOR in detecting extended galaxies, we consider mock images of six CASTOR pointings, taken from independent, non-overlapping regions of the input 2 deg2 catalogue. This allows us to account for some cosmic variance. We extract and then match all detected galaxies within each of these six images, for each filter and survey exposure time.
@@
The Ultra-Deep survey with the largest exposure times clearly detects the most galaxies, with the largest detection fractions out to the faintest magnitudes. As shown in Table 2, the 75% completeness limits for this Ultra-Deep survey are 28.3–29.2 mag, with galaxies as faint as mag detected. The filters are well matched in depth within this Ultra-Deep survey. The Deep survey will reach around 1.3–1.6 magnitudes brighter than the Ultra-Deep survey, with a factor of 10 less exposure time.
The significantly shallower Wide survey will be 75% complete between 22.8–25.0 mag, approximately 4–5 mag fainter than the Ultra-Deep survey, and will reach depths of up to 27.5 mag. In this survey the decreased sensitivity of the UV-split and u-split filters is clear, resulting in a decreased completeness limit of up to 2 mag compared to their un-split counterparts. These narrower filters will play a key role in understanding the properties of the sources detected in CASTOR, providing further data points for SED fitting, as well as being combined with the un-split filter images to create even deeper detection images, however in general the UV, u and g filters will be the optimal filters for source detection in the Wide survey.
To further compare the performance of each survey, we list the number of galaxies detected in each CASTOR FOV in Table 3, as well as an estimation of the total number of galaxies in the full surveys by extrapolating this value across the survey areas. In one pointing, the Ultra-Deep survey is expected to detect an order of magnitude more galaxies than the Wide survey, of order 100,000 galaxies compared to . However, with its large survey area of 2227 deg2, this will correspond to detected galaxies across the whole Wide survey. The much deeper Ultra-Deep survey’s small volume of only 4 CASTOR pointings (1 deg2), will detect galaxies, albeit down to much fainter magnitudes. The intermediate Deep survey, covering 83 deg2, will detect galaxies.
@@ 5.3 Completeness relative to galaxy properties
We now investigate the detectability of galaxies in each CASTOR survey relative to the intrinsic galaxy properties from the input catalogue, for galaxies brighter than mag. In Figure 6 we show the fraction of input galaxies detected as a function of stellar mass for each survey. This is shown for each of the filters, and split into redshift bins. The detection fraction generally increases with galaxy stellar mass, with a plateau at large mass depending on the filter and exposure time. Generally, we expect the Wide survey to detect very few galaxies with . The Deep survey may detect a small fraction of galaxies down to , while the Ultra-Deep survey may detect galaxies with at rates of , as well as a small fraction of galaxies down to .
In Figure 7 we show the equivalent figure of the fraction of galaxies detected as a function of galaxy disk radius. Galaxies with very small disk sizes of kpc have much lower rates of successful detection. The detection fractions also trend downwards for larger galaxies, particularly in the Wide survey, which is likely due to their lower surface brightnesses making them more difficult to observe. The average detection fraction of disks with kpc depends strongly on exposure time and filter, with the Wide survey having detection fractions of , up to with the Ultra-Deep survey.
@@ 5.4 Magnitude recovery
Finally we consider how the input and recovered magnitudes compare. In Figure 9 we show the extracted ISO magnitudes from Source Extractor, compared with the input magnitude. We note that throughout this paper we have otherwise considered only the input theoretical magnitude.
We find that in the Wide survey, 95% of all galaxies have an extracted magnitude within 3.3%, 4.1%, 2.9%, 3.2% and 2.9% of the input magnitude, for the UV, UV-split, u, u-split, and g filters respectively. These extracted magnitudes give accurate measurements of the input magnitudes, with a median bias of 0.6% above the input magnitudes. These are slightly improved in the Deep and Ultra-Deep surveys, to 2.0%, 2.2%, 1.9%, 2.0% and 2.0% for Deep, and 1.7%, 1.8%, 1.6%, 1.7% and 1.7% for Ultra-Deep. The extracted magnitudes from both the Deep and Ultra-Deep surveys have a median bias of 0.4% above the input magnitudes. The low number of galaxies with larger discrepancies between the input and output magnitudes are very faint.
The extracted magnitudes are both accurate and precise, with median bias , and with 95% of galaxies having magnitudes correct to within , for all filters in the three surveys. Overall, the CASTOR surveys can accurately measure galaxy magnitudes.
@@ 6.1 Synergy with Euclid and Roman
CASTOR’s wide-field survey capabilities and UV coverage will play an indispensable role in the new era of wide-field near-infrared extragalactic surveys, alongside Roman (Spergel et al., 2013, 2015) and Euclid (Laureijs et al., 2011; Racca et al., 2016). These space-based observatories have fields of view of the order of hundred times bigger than Hubble’s (0.54 deg2 for Euclid and 0.28 deg2 for Roman) and possess angular resolutions comparable to that of Hubble (with a mirror diameter of 1.2 m for Euclid and 2.4 m for Roman), similar to those of CASTOR (0.25 deg2 and m). Roman and Euclid focus on the near-infrared and optical, covering wavelengths up to m. However, as these will provide no wavelength coverage below Å, CASTOR will play a key role with its wavelength coverage that extends beyond the optical to the near UV. Together, these three observatories are set to transform galaxy and cosmological surveys. Harnessing the synergy between CASTOR and other observatories will be the key to maximizing the scientific return across all observatories.
@@ 7 Conclusions
The Cosmological Advanced Survey Telescope for Optical and UV Research (CASTOR) is a planned Canadian-led space telescope designed to study the Universe in the UV and optical. A core component of the planning for CASTOR is the preparation of a suite of mission planning and science simulation tools, the Finding Optical Requirements and Exposure times for CASTOR (FORECASTOR) project. In this work we present the publicly available CASTOR image simulator222https://github.com/CASTOR-telescope/ImageSimulator, a generator which simulates extragalactic imaging for CASTOR observation planning. This generator is built upon the framework of the GalSim Python package, and creates images containing sources from simulated light-cones from the Santa Cruz Semi-Analytic Model.
@@
Three main extragalactic surveys planned for the mission are the Wide, Deep, and Ultra-Deep Legacy Surveys, spanning 2227, 83 and 1 deg2, respectively. These surveys will reach total exposure times of 1000, 18,000 and 180,000s respectively in the UV, UV-split, u and u-split filters, with twice the exposure time in the g filter. In this paper we present mock images of each of these surveys, and investigate the detectability of galaxies at these planned exposure times. These images are publicly available333https://doi.org/10.11570/24.0003, to enable further investigations of the extragalactic science achievable with CASTOR. The resolution of CASTOR will result in crisp UV images, which will significantly out-perform GALEX with its that resulted in significant galaxy blending. This is also a higher resolution than for Rubin, which is limited to due to atmospheric seeing.
@@
We extract sources from the images and compare those to the input source catalogue, in order to estimate the completeness of each CASTOR survey. For point sources, we predict that the Wide survey will be 75% complete down to 26.2–27.2 mag, with the Deep survey extending to 28.2–28.8 mag, and the Ultra-Deep survey reaching to 29.5–30.2 mag. The five filters, UV, UV-split, u, u-split, and g, are generally well matched in depth for each survey. For galaxies, we expect that the faintest detectable galaxy discovered in the Wide survey will have mag, with Deep reaching to mag and Ultra-Deep discovering some faint sources out to mag. The recovered magnitudes are in good agreement with those from the input catalogue, with median bias , and with 95% of galaxies having magnitudes correct to within , for all filters in the three surveys. For the completeness relative to stellar mass, we expect the Wide survey to detect very few galaxies with . The Deep survey may detect galaxies down to , while the Ultra-Deep survey may detect galaxies down to . Galaxies with disk radii of kpc will be the most difficult to detect.
@@
For a single 0.25 deg2 pointing, the Wide survey is expected to detect of order galaxies, with galaxies seen in one pointing of the Deep and Ultra-Deep surveys. Extending these counts to the full survey areas, we expect the Wide survey to detect galaxies, with galaxies in the Deep survey and in the Ultra-Deep survey.
@@
Acknowledgements
We sincerely thank the referee for their supportive and helpful feedback on this project.
MAM acknowledges support by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20240752PRD1.
This research used the Canadian Advanced Network For Astronomy Research (CANFAR) operated in partnership by the Canadian Astronomy Data Centre and The Digital Research Alliance of Canada with support from the National Research Council of Canada, the Canadian Space Agency, CANARIE, and the Canadian Foundation for Innovation.
This paper made use of Source Extractor (Bertin & Arnouts, 1996), and Python packages and software AstroPy (Astropy Collaboration et al., 2013, 2018), GalSim (Rowe et al., 2015), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011), Pandas (Pandas Development Team, 2020), Photutils (Bradley et al., 2018), SciPy (Virtanen et al., 2020), and Seaborn (Waskom, 2021).
Data Availability
The mock CASTOR survey images generated in this work are available on CANFAR, at https://doi.org/10.11570/24.0003. The CASTOR image simulator, including the currently estimated CASTOR filter bandpasses and PSFs, is available at https://github.com/CASTOR-telescope/ImageSimulator. The simulated lightcones used in this work are not publicly available, but will be shared on reasonable request to the corresponding author.